OpenStudy (anonymous):

5/9 divided by 3 1/3

OpenStudy (anonymous):

1/6

OpenStudy (anonymous):

Does it look like this \[\frac{ \frac{ 5 }{ 9 } }{ 3\frac{ 1 }{ 3 } }\]?

OpenStudy (anonymous):

the answer is 1/6. hes right

OpenStudy (anonymous):

Did you understand how to get that answer though?

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

No how?

OpenStudy (anonymous):

You turn the 3 1/3 into an improper fraction so it would turn into 10/3. does that make sense? and instead of dividing the two fractions you would multiply but using the reciprocal of a fraction

OpenStudy (anonymous):

Thank u. Soooo much

OpenStudy (anonymous):

\[\frac{ \frac{ 5 }{ 9 } }{ \frac{ 10 }{ 3 } }\] becomes \[\frac{ 5 }{ 9 } * \frac{ 3 }{ 10 }\] then you simplify

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

Thank you this is graded i needed this

OpenStudy (anonymous):

I hope you get a good mark! goodluck!